How much was borrowed is a simple interest rate of 2.10% was offered payable in 4 months with interest of Php 35.00?

1
Expert's answer
2021-12-14T12:02:20-0500

"Principle 

= \frac{Simple\space interest }{ (interest \times time)}\\

= \frac{35 }{ (2.10\%\times \frac{4}{12})}\\

=5000.00\\


Amount\space borrowed =

5000.00"